# Enhancing Object Detection Accuracy in Underwater Sonar Images through Deep Learning-based Denoising

### v1

@@ III-A Sonar Images
Typically, sonar images consist of seven main components: the sonar pulse source, surface reflections, water column clutter, seabed reflections, water column, target objects, and shadows. The process of image formation by sonar equipment generally includes the following steps: the transducer array emits a sonar pulse and listens for returning echoes. The returned sound waves are converted from acoustic energy into electrical energy, which is transmitted via a towed cable to the recording and display unit onboard the surface vessel. At the recording and display unit, the signal is processed into the desired format and then sent to monitors and recorders, resulting in a recorded image. The main characteristics of sonar images are as follows:
Low Resolution. Due to the long wavelength of sound waves, sonar images typically have lower resolution, resulting in blurry details.
Noise Interference. Sonar images are often affected by various noise sources, such as speckle noise and multipath effects, creating noticeable noise patterns.
Low Signal-to-Noise Ratio (SNR). Given the complex underwater environment, sonar imaging often has a low SNR, making the contours of objects blurry and harder to identify.
Depth Information. Sonar images provide valuable depth information about objects, facilitating the identification and localization of targets in a 3D underwater environment.
Fig. 1 presents two images: the top one is a side-scan sonar image, while the bottom one is a forward-looking sonar image. In the side-scan sonar image, the black area in the center represents sonar shadows (as the detector emits sound waves to both sides, the area directly beneath receives no echoes, forming a black region). The bright yellow areas on both sides indicate the returned sonar signals, revealing details of the seabed terrain, such as sand dunes or other geological features. Since side-scan sonar scans the seabed from the sides, it provides a wide coverage area, making it suitable for underwater terrain surveying and target detection. The forward-looking sonar image contains numerous radial streaks, which may be noise caused by underwater target reflections or sound wave interference. The primary purpose of forward-looking sonar is to detect obstacles ahead of underwater vehicles or divers, providing real-time navigation information.
III-B Noise in sonar images
Sonar images are affected by a variety of noise types, including Gaussian noise, Poisson noise, speckle noise, and other environmental noises. Each type of noise impacts the quality of sonar images and the accuracy of object detection in distinct ways.
Gaussian Noise. Gaussian noise typically arises from sensor instability and environmental factors, and it is randomly distributed throughout the image. This type of noise significantly affects the overall image quality, especially under low-light conditions. Due to its randomness, Gaussian noise is evenly spread across the image, blurring details of the target and making precise object recognition and detection more challenging.
Poisson Noise.Poisson noise occurs primarily in low-light or low-signal scenarios, with its intensity proportional to the signal strength. It is especially noticeable in underexposed regions of the image. In sonar images, Poisson noise can reduce the visibility of the signal, blurring target edges and making image processing and object detection more difficult.
Speckle Noise. Speckle noise is a common noise type in sonar images, generally caused by sound wave reflection and scattering. It appears as granular interference, giving the image an uneven texture and causing blurred target boundaries. Speckle noise significantly impacts image detail, especially in high-resolution images, and can impair the accuracy of object detection.
Other Noises. Apart from the primary noise types, sonar images may also contain other noises, such as multipath reflection noise and environmental noise. Multipath reflection noise arises from variations in the reflection paths of sound waves, potentially causing image blurring or false targets. Environmental noise may stem from underwater biological activity, ship vibrations, and other external interferences, adding to the complexity of noise in sonar images. These various types of noise present significant challenges for object detection in underwater sonar images.
III-C Typical Object Detection Algorithms
The development of object detection algorithms has gone through several important stages, gradually transitioning from traditional detection methods to real-time detection technologies based on deep learning. YOLO Series. YOLO [24] was proposed by Joseph et al. in 2015 and is the first one-stage detector in the deep learning era. YOLO is extremely fast, following a completely different paradigm from two-stage detectors. It applies a single neural network to the full image, dividing the image into regions and predicting bounding boxes and probabilities for each region simultaneously. Despite its significant improvement in detection speed, YOLO suffers from a drop in localization accuracy, especially for small objects. Subsequent versions of YOLO (such as YOLOv4 [25] and YOLOv7 [26]) focus on addressing this issue. YOLOv7 introduces optimized structures like dynamic label assignment and model structure reparameterization, outperforming most existing object detectors in speed and accuracy. YOLOv5 [44] builds on YOLOv4 by optimizing model architecture and hyperparameters, making training and deployment more streamlined while enhancing speed and accuracy. In addition, YOLOX [45], a newer member of the YOLO family, incorporates anchor-free mechanisms and decoupled heads, further boosting its detection capabilities while maintaining high efficiency. SSD Series. SSD [27] was proposed by Liu et al. in 2015. Its main contribution is the introduction of multi-reference and multi-resolution detection techniques, significantly improving the detection accuracy of a one-stage detector, especially for small objects. SSD detects objects of different scales on different layers of the network, whereas previous detectors only ran detection on their top layers. R-CNN Series. The idea behind R-CNN [28] is straightforward. It begins with the extraction of a set of object proposals (candidate boxes) using selective search. Each proposal is then rescaled to a fixed size and fed into a CNN model pretrained on ImageNet (e.g., AlexNet[43]) to extract features. Finally, linear SVM classifiers are used to predict the presence of an object within each region and recognize object categories. However, the drawbacks are evident: redundant feature computations on a large number of overlapping proposals (over 2000 boxes from one image) lead to an extremely slow detection speed. Later in the same year, SPPNet [29] was proposed to solve this problem. In 2015, Girshick proposed Fast R-CNN [30], a further improvement of R-CNN and SPPNet. Fast R-CNN enables simultaneous training of a detector and a bounding box regressor under the same network configurations. Although Fast R-CNN successfully integrates the advantages of R-CNN and SPPNet, its detection speed is still limited by proposal detection. Ren et al. proposed Faster R-CNN [31] shortly after Fast R-CNN. Faster R-CNN is the first near-real-time deep learning detector. Its main contribution is the introduction of a Region Proposal Network (RPN), enabling nearly cost-free region proposals.
IV Methodology
The variety and abundance of natural image datasets have fueled the development of advanced denoising methods, especially self-supervised algorithms. These algorithms do not rely on large-scale labeled datasets, making them well-suited for noise removal tasks in scenarios with scarce labeled data. Underwater sonar imaging faces significant challenges due to the diverse sonar devices, complex underwater environments, and varying noise patterns. These noise types, such as speckle noise, gaussian Noise, and poisson noise, significantly degrade the quality of sonar images, complicating subsequent analysis and object detection tasks. Unlike natural image datasets, which often include clean or minimally noisy images, obtaining noise-free sonar images is exceptionally difficult. This limitation necessitates the adaptive application of existing denoising methods. Consequently, these algorithms require tailored evaluation and adjustment to verify their effectiveness in sonar imaging. We designed a method to verify whether applying denoising algorithms originally developed for optical images to sonar image datasets can improve object detection accuracy and whether performing multi-frame denoising on images processed by different algorithms can further enhance detection precision.
Fig. 2 presents the proposed method designed to validate the above questions. This section introduces the proposed methodology for validating the impact of denoising techniques and fusion strategies on underwater sonar image object detection. The workflow is divided into three major steps: baseline object detection, object detection on denoised datasets, and object detection on multi-frame denoised datasets. Each step is detailed below to explain the overall process and its rationale.
First, multiple original underwater sonar image datasets (Dataset1, Dataset2, …, DatasetX) are processed using several object detection algorithms (Object Detection Algorithm1, Algorithm2, …, AlgorithmX) to establish a baseline for comparison. Next, the original datasets are denoised using state-of-the-art self-supervised denoising algorithms (Self-Supervised Denoising Algorithm1, Algorithm2, …, AlgorithmX), generating denoised datasets (Denoised Dataset1, Denoised Dataset2, …, Denoised DatasetX). These denoised datasets are subsequently evaluated using the same object detection algorithms. Finally, multi-frame image denoising techniques are employed to further denoise the denoised dataset, generating multi-frame denoised datasets (Multi-frame Denoised Dataset 1, Multi-frame Denoised Dataset 2, …, Multi-frame Denoised Dataset X). Object detection is then performed on these multi-frame denoised datasets to assess potential improvements in detection accuracy and robustness. This method systematically explores how denoising and multi-frame denoising can enhance detection outcomes while preserving critical information in sonar images.
IV-A Object Detection on Original Datasets
The first step involves processing original underwater sonar image datasets to establish a baseline for object detection performance. To establish a performance baseline, we collect sonar image datasets using various sonar devices and in multiple underwater scenarios. These datasets represent the diversity of noise types and environmental conditions encountered in real-world applications. The effectiveness of object detection algorithms is evaluated on these datasets, enabling us to identify the challenges posed by different noise patterns and device characteristics.
IV-B Object Detection on Denoised Datasets
Given that underwater sonar images often lack clean, noise-free ground truths, self-supervised denoising algorithms are utilized. These algorithms (e.g., Self-Supervised Denoising Algorithm1, Algorithm2, …, AlgorithmX) are specifically chosen for their ability to learn noise patterns without relying on clean ground-truth images. The denoised datasets are subsequently processed using the same object detection algorithms applied in the baseline step. Comparing detection results on denoised datasets with the baseline allows us to quantify each denoising algorithm’s effectiveness. This step examines the impact of noise reduction on object detection performance and evaluates whether certain algorithms perform better for specific sonar noise types.
IV-C Enhancing Datasets with Multi-Frame Denoising Models
To further improve object detection performance, we consider the denoised datasets as multiple frames of the same scene captured in sequence. This approach is based on the premise that applying different denoising algorithms to the same image produces multiple versions, each with unique strengths in noise suppression and detail preservation, resembling observations of the same scene under varying conditions. Therefore, we utilize these denoised datasets as training data to train a deep learning-based multi-frame denoising model, enabling it to extract stable features from diverse denoised versions while enhancing noise adaptability.
Once trained, this model is applied to the original sonar dataset for denoising, aiming to reduce noise while preserving critical information essential for object detection. Finally, object detection is performed on the model-denoised dataset to evaluate its impact on detection accuracy and robustness. This multi-frame denoising strategy not only leverages the complementary strengths of different denoising algorithms but also enhances the model’s generalization ability to complex underwater sonar noise, providing a more effective data preprocessing solution for underwater object detection.
By systematically comparing the results from original, denoised, and multi-frame denoised datasets, this methodology provides valuable insights into the interplay between denoising, multi-frame denoising, and object detection performance in underwater sonar imagery.
V Experiments

### v2

@@ 2.1 Sonar Images
A typical sonar image consists of seven key components: the sonar pulse source, surface reflections, water column clutter, seabed reflections, the water column itself, target objects, and associated shadows. The image formation process begins with the transducer array emitting a sonar pulse and subsequently capturing the returning echoes. These acoustic signals are converted into electrical energy and transmitted via a towed cable to the surface vessel’s recording and display unit, where they are processed into interpretable formats and rendered on monitors or stored for further analysis.
Sonar images exhibit several distinctive characteristics. First, their resolution is inherently limited due to the long wavelengths of sound waves, resulting in comparatively blurred details. Second, they are prone to noise interference, including speckle noise and multipath effects, which can introduce artifacts into the imagery. Third, the complex underwater environment often leads to a low signal-to-noise ratio (SNR), further obscuring object contours and complicating identification. Despite these challenges, sonar images provide invaluable depth information, enabling effective target localization and navigation in three-dimensional underwater spaces.
2.2 Noise in Sonar Images
The interpretation of sonar images is fundamentally constrained by various noise sources that degrade image quality and complicate object detection. These noise artifacts originate from different physical mechanisms in the underwater environment and sensor systems. The primary noise types affecting sonar images include:
1.
Gaussian Noise: This additive noise follows a normal distribution and primarily stems from electronic sensor noise and environmental interference. It manifests as random intensity fluctuations uniformly distributed across the image, particularly problematic in low-signal conditions where it obscures fine details and reduces contrast.
2.
Poisson Noise: Also known as shot noise, this signal-dependent noise arises from the discrete nature of acoustic wave detection. Its variance scales with signal intensity, making it particularly noticeable in low-intensity regions of the image where it introduces grainy artifacts and distorts weak signals.
3.
Speckle Noise: As a multiplicative noise inherent to coherent imaging systems, speckle noise results from interference patterns created by sound wave scattering. It appears as granular texture throughout the image, significantly blurring object boundaries and reducing effective resolution.
4.
Multipath Interference: Caused by sound waves reflecting off multiple surfaces before detection, this noise creates ghost images and false targets. The complex underwater topography and variable acoustic impedance contribute to this challenging interference pattern.
5.
Environmental Noise: This category encompasses various ambient disturbances including biological activity, water turbulence, and man-made sources like vessel operations. These noises are typically non-stationary and spatially varying, requiring adaptive processing techniques.
The combined effect of these noise sources presents significant challenges for automated object detection and classification in sonar imagery. Effective noise suppression and image enhancement algorithms must account for these diverse noise characteristics to improve the reliability of underwater imaging systems.
3 Related Work
3.1 Underwater Sonar Image Object Detection
Object detection in underwater sonar imagery encompasses traditional methods, convolutional neural network (CNN)-based approaches, and other machine learning techniques.
Traditional sonar detection relies on pixel intensities, grayscale thresholds, and target priors. For example, [20] employs spatial frequency domain analysis and frame differencing for robust detection in underwater videos. Similarly, [21] uses OS-CFAR for noise-resilient target segmentation in side-scan sonar, validated on real data. Additionally, [22] applies symbolic pattern analysis for real-time mine detection in sidescan sonar, while [23] explores change detection to identify new seabed objects.
CNN-based methods excel in sonar imagery due to robust feature extraction[24, 25, 26, 27]. YOLO-based models include YOLOv3-DPFIN [28], which enhances real-time detection with dual-path feature fusion, and an optimized YOLOv5 [29] using transfer learning and CoordConv for improved accuracy. YOLOv7 variants, YOLOv7C [30] with attention mechanisms and Multi-GnBlock, and a Swin-Transformer-enhanced model [31] with a feature scale factor, boost performance in complex sonar backgrounds. Other CNN approaches include a lightweight FLSD-Net [32] for small object detection, UFIDNet [33] with speckle noise reduction, EPL-UFLSID [34] using pseudo labels, and MB-CEDN [35] for multi-target detection in circular-scan sonar, all improving accuracy and efficiency. Additionally, [36] outperforms template matching in forward-looking sonar with fewer parameters.
Non-CNN methods include a Viola-Jones-based cascade classifier [37] for in-situ learning and low false alarms in complex terrains, and a clustering-segmentation approach [38] for efficient, low-error detection without extensive training.
Traditional methods utilize image processing and handcrafted features for rule-based detection, while CNN-based methods leverage end-to-end feature learning for enhanced robustness. Other machine learning approaches combine manual features with classifiers, offering flexibility but limited by feature quality.
3.2 Denoising
Image denoising is critical in image processing, particularly for noisy underwater sonar imagery. Deep learning, especially convolutional neural network (CNN)-based methods, dominates denoising due to robust performance. RED-Net [39] uses skip-layer connections for enhanced denoising, while DnCNN [40] handles varying noise intensities for blind denoising. FFDNet [41] incorporates a noise level map to improve inference speed, and CBDNet [42] employs realistic noise models for real-world applications. MWDCNN [43] integrates wavelet transforms and dynamic convolution for balanced performance, whereas CTNet [44] combines serial, parallel, and residual blocks with Transformer mechanisms to address complex noise in large datasets.
Unsupervised and self-supervised methods enable denoising without clean images. Noise2Noise [45] trains on noisy image pairs, recovering clean images effectively. Noise2Self [46] and Noise2Void [47] exploit noise statistics for self-supervised denoising, while Noisy-As-Clean [48] treats noisy images as clean to reduce prior discrepancies. Neighbor2Neighbor [49] and Blind2Unblind [50] further advance self-supervised denoising with efficient training.
For domain-specific tasks, LIDnet [51] integrates denoising with lesion detection in medical imaging, using region-specific loss and collaborative training. DN-DETR [52] enhances DETR’s convergence via denoising strategies. In sonar imagery, UFIDNet [33] combines speckle noise reduction with feature selection for accurate forward-looking sonar detection, while NACA [53] leverages non-local spatial information and adaptive algorithms for robust denoising and detection. EPL-UFLSID[34] method proposes a Gaussian Mixture Model-based Deep Image Prior network for generating denoised sonar images, utilizing the GMM distribution as its input. UGIF-Net[54] enhances underwater images using a multicolor space-guided module and dense attention block. HCLR-Net[55] enhances underwater images using a hybrid contrastive learning regularization strategy with non-paired data and local patch perturbations,
Underwater sonar denoising remains nascent, with future work needed to address complex noise types and adapt optical image denoising techniques for sonar applications.
4 Methodology
Our methodology is structured in three parts: (1) presentation of used four representative object detection algorithms, (2) description of used nine denoising models, and (3) comprehensive explanation of our proposed M2SDF framework method. The overall methodology workflow is illustrated in Fig. 2.
4.1 Baseline Object Detection Architecture Selection
To establish a comprehensive evaluation framework for denoising-enhanced sonar object detection, we selected four representative detection algorithms spanning distinct architectural paradigms: the lightweight SSDMobileNetV2 [56], the two-stage Faster R-CNN [57], the single-stage SSD300 [58], and the anchor-free YOLOX [59]. This curated selection enables systematic analysis of denoising effects across efficiency-accuracy tradeoffs.
SSDMobileNetV2: We include SSDMobileNetV2 due to its lightweight architecture and computational efficiency, making it well-suited for real-time applications on resource-constrained devices. Its backbone, MobileNetV2 [60], employs two key innovations: (1) inverted residual blocks with linear bottlenecks and (2) depthwise separable convolutions. The inverted residual structure facilitates gradient propagation via shortcut connections while maintaining low computational overhead, whereas linear bottlenecks preserve information by avoiding nonlinearities in narrow layers. These design choices enable SSDMobileNetV2 to balance accuracy and efficiency effectively, allowing us to evaluate denoising’s impact on lightweight models optimized for mobile or embedded deployment.
Faster R-CNN: Faster R-CNN is selected for its high localization and classification accuracy, particularly in complex environments. Its architecture consists of a Region Proposal Network (RPN) for generating candidate regions, followed by a second-stage classifier and regressor for refined detection. Typically paired with deep backbones like ResNet, Faster R-CNN excels in scenarios with small, dense, or partially obscured objects—common challenges in sonar imagery. Although computationally intensive, its precision provides a valuable benchmark for assessing denoising’s effect on detection performance in high-complexity scenes where precise object delineation is critical.
SSD300: SSD300 directly predicts object categories and bounding boxes from multi-scale feature maps, eliminating the need for a separate region proposal step. Operating on 300×300 inputs, it leverages multi-scale convolutional layers to handle objects of varying sizes—an advantage for sonar imagery, where targets exhibit significant scale variations and noise interference. Its real-time performance and multi-scale adaptability make it ideal for studying denoising’s influence on mid-range models that prioritize both efficiency and reasonable accuracy.
YOLOX: YOLOX represents a state-of-the-art advancement in one-stage detection, addressing limitations of earlier YOLO variants. It introduces an anchor-free design, a decoupled head architecture, and SimOTA label assignment to improve accuracy and training stability. By leveraging modern backbones (e.g., CSPDarkNet or EfficientNet [61]) and incorporating speed optimizations, YOLOX achieves robust performance in both real-time and high-precision tasks. Its resilience in complex detection scenarios makes it a strong baseline for evaluating denoising’s impact on high-performance models employing contemporary design principles.
4.2 Single-Frame Denoising Algorithm Analysis
We select three state-of-the-art deep learning-based denoising algorithms, each offering unique advantages and configurable parameters to address diverse noise types. By adjusting their key settings, we derive nine distinct denoising variants. Below, we detail these algorithms and our configuration strategy.
4.2.1 Algorithm 1: Neighbor2Neighbor
Neighbor2Neighbor [49] is a self-supervised image denoising framework that operates exclusively on single noisy images, eliminating the need for clean reference images or multiple observations of the same scene. This characteristic makes it especially applicable to challenging domains like underwater sonar image processing, where obtaining clean ground-truth data is impractical. The method introduces two key innovations:
1.
Random Neighbor Sub-sampling: Rather than depending on noisy-clean image pairs, Neighbor2Neighbor generates pseudo training pairs directly from a single noisy input. By randomly extracting two adjacent sub-regions from the image—which exhibit statistical similarity due to spatial redundancy—the method creates input-target pairs for training. This approach capitalizes on the inherent local self-similarity present in natural and sonar imagery.
2.
Regularized Objective Function: To address the inherent discrepancy between the sub-sampled patches (since their underlying clean versions are not identical), the algorithm incorporates a theoretical correction mechanism through a regularization term. This component mitigates the mismatch arising from the pseudo pairs, enhancing denoising accuracy while preventing excessive smoothing.
Furthermore, Neighbor2Neighbor is architecture-agnostic, enabling seamless integration with state-of-the-art supervised denoising networks. Its noise-model-free design enhances robustness for real-world scenarios, such as sonar imaging, where noise patterns are often non-Gaussian and complex.
We adopt Neighbor2Neighbor as a representative self-supervised denoising method to assess the influence of denoising on sonar-based object detection. Its independence from clean images aligns perfectly with sonar datasets, which typically lack noise-free references. Additionally, its ability to deliver high-quality denoised outputs using only noisy inputs ensures our evaluation reflects real-world operational constraints.
4.2.2 Algorithm 2: Blind2Unblind
Blind2Unblind [50] addresses the critical shortcomings of blind-spot-based self-supervised approaches, which typically incur substantial information loss from input masking or constrained receptive fields. The method introduces two innovative components to mitigate these issues:
1.
Global-Aware Mask Mapper:The algorithm divides the input image into small grid cells (e.g., 2×2 patches) and systematically masks individual pixels to construct a masked volume—a stack of variably masked images. A denoising network processes this volume, and the mapper intelligently aggregates predictions from all masked regions to reconstruct the final output. This mechanism enables holistic image understanding while significantly improving training convergence.
2.
egularized Re-Visible Loss: To bridge the gap between blind-spot training and full-image inference, the method proposes a novel loss function combining a regularization term and a dynamically weighted visible term. This formulation prevents degenerate solutions (e.g., identity mapping) and stabilizes optimization. Notably, the loss is backbone-agnostic, ensuring compatibility with diverse denoising architectures.
Blind2Unblind is suited for sonar data, where noise (e.g., speckle, multipath interference, and non-Gaussian ambient noise) complicates traditional denoising, and clean reference images are rarely available. By leveraging global context and operating entirely self-supervised, the method offers a robust solution for real-world sonar applications.
4.2.3 Algorithm 3: DSPNet
Speckle noise presents a fundamental challenge in sonar image processing due to its multiplicative nature. Traditional denoising methods often fail to adequately address this problem as they typically assume additive noise models. The DSPNet framework [62] overcomes this limitation through a novel logarithmic transformation approach that effectively converts multiplicative speckle noise into an additive model.
The architecture employs two specialized subnetworks working in the logarithmic domain. First, the noise estimation network (Log-NENet) utilizes a U-Net structure with attention mechanisms to predict spatially-varying noise levels. Second, the denoising network (Log-DNNet) incorporates a feature pyramid network with atrous spatial pyramid pooling (ASPP) modules, enabling multi-scale feature extraction while preserving fine image details. This dual-network approach achieves superior performance by explicitly modeling the noise characteristics while maintaining structural integrity.
A key innovation of DSPNet lies in its hybrid loss function, which combines multi-scale structural similarity (MS-SSIM) with traditional L1 loss. This combination ensures both perceptual quality and pixel-level accuracy in the denoised output.
For sonar applications, DSPNet offers particular advantages. Its logarithmic processing naturally aligns with the physics of underwater acoustic imaging, where speckle noise follows multiplicative Rayleigh distributions. Furthermore, the network’s ability to operate without clean reference images makes it practical for real-world deployment scenarios where ground truth data is unavailable.
4.2.4 Denoising Model Configuration
To comprehensively evaluate denoising performance under various noise conditions, we established a systematic training protocol encompassing nine specialized models, includes:
We implemented four variants each for both Blind2Unblind and Neighbor2Neighbor by training under distinct noise regimes:
1.
Fixed-intensity Gaussian noise
2.
Variable Gaussian noise ()
3.
Fixed-parameter Poisson noise
4.
Variable Poisson noise
This configuration enables comparative analysis of: 1) Model robustness to noise level variations; 2) Algorithmic performance differences between fixed and variable noise conditions; 3) Generalization capability across noise types.
The ninth model employs DSPNet, specifically designed for multiplicative speckle noise reduction through its logarithmic domain processing. This provides critical comparison with the additive noise models, particularly relevant for sonar imaging where speckle noise dominates.
4.3 Mutually-supervised Multi-source Denoising Fusion
To enhance underwater sonar target detection, we propose M2SDF, a novel framework that integrates multi-source heterogeneous denoisers to synergize their complementary strengths. Inspired by multi-frame denoising principles [63], M2SDF employs cross-referenced pixel-wise supervision among denoised outputs from different denoisers. This mutual supervision enables the model to learn more stable and generalizable features by leveraging inter-denoiser consistency, thereby improving robustness in complex noise environments.
We first apply a set of single-frame denoising methods to each noisy image from the original dataset D. Each method generates a denoised variant:
These denoised variants of the same image are grouped to form a multi-source sequence:
This sequence is then used as the input to M2SDF with parameters , producing a final denoised output:
For each image, the corresponding noisy frames are randomly shuffled before each training iteration to form new supervision pairs. This randomization introduces dynamic supervisory directions, allowing the model to treat each frame both as input and as a supervisory target in different iterations. At each training step, two disjoint subsets of frames are formed, and the loss is computed between these paired groups. With sufficient training steps, this strategy achieves balanced and diversified mutual supervision (as illustrated in Fig.3). The randomized re-pairing of frames at each iteration effectively enriches the diversity of training signals without increasing computational overhead.
During the model training phase, the model processes noisy inputs in an iterative manner. To enable dynamic and balanced supervision across frames, we introduce a random shuffling strategy at each iteration to construct paired data samples. Let s denote the current iteration step, and define a frame index sequence . Before the -th update, this sequence is randomly divided into two disjoint subsets , and , for each training sample , where . Based on these subsets, the corresponding noisy frame groups are formed as:
Then, the model parameters are updated at each iteration using the following rule:
Where:
1.
is the model parameters at the -th iteration;
2.
is the learning rate;
3.
The constraint ensures that the paired samples originate from the same data instance.
This randomized re-pairing mechanism ensures that every noisy frame in a sequence has an equal opportunity to act as either the supervision target or the denoising input over time. In effect, this method gradually builds implicit multi-directional supervision relationships across all frames without requiring clean labels or reference images. As the number of iterations increases, the statistical uniformity of frame usage improves, fostering a diverse and balanced training process that encourages generalization while maintaining computational efficiency.
Our approach extends self-supervised learning by jointly optimizing outputs of different single-image denoisers (Neighbor2Neighbor, Blind2Unblind, etc.). Unlike traditional methods requiring clean references or true multi-frame captures, our framework capitalizes on intrinsic discrepancies among different denoiser outputs - including noise suppression levels, texture preservation characteristics, and artifact patterns. This enables mutual supervision through consensus learning in stable regions and discrepancy-based correction in noisy areas.
5 Experiments